4.8 and 4.9

A sample of 40 sales receipts from a university bookstore has x= $121 and s-$102. Use these values to perform a test of Ho: ?-125 against Hi: ? < 125 with ? 05. 4.8 4.9 Use the values in Exercise 4.8 to test Ho: ?:: 125 versus H1 : ? 125 with ? = .05.
Answer #1

4.8

t-statistic can be calculated by

t=(121-125)/(10.2/sqrt(40)) = -2.48

since it is a one-tailed hypothesis test

critical t-statistic = t(0.05,39) = -1.685

Icalculated t-statisticI>Icritical t-statisticI

There are not enough evidence to accept the null hypothesis.

Alternate hypotheis is accepted.

4.9

calculated t-statistic=-2.48

since it is a two-tailed hypothesis test

critical t-statistic = t(0.025,39) =-2.022

Icalculated t-statisticI>Icritical t-statisticI

There are not enough evidence to accept the null hypothesis.

Alternate hypotheis is accepted.